# GeoGPT.RAG Technical Report

### v1

@@ 1 Introduction
  • LLM-based synthetic data generation pipeline for geoscience retrieval. To address the lack of high-quality labeled data in the geoscience domain, we design a novel synthetic data generation pipeline powered by LLaMA3-70B. This pipeline produces realistic and diverse training data by abstracting and reformatting scholarly geoscience articles, enabling robust training of retrieval and reranking models under data-scarce conditions.
@@
  • GeoQA: a new benchmark for geoscience information retrieval. We create and release GeoQA, the first large-scale benchmark designed for evaluating geoscience retrieval systems. It contains OA paper-based domain-specific documents and 938 expert-reviewed query-document pairs, providing a standardized framework for measuring performance and guiding future research in this area.
@@ 3 RAG Evaluation
We evaluated GeoGPT’s retrieval and generation capabilities using a custom test set generated with the Test Set Generation module from RAGAS (Es et al., 2023). This test set, named GeoQA, comprises 1,000 questions derived from publicly available geoscience papers. We used gpt-4o as the underlying LLM for both question-answer (QA) pair generation and answer evaluation, and text-embedding-3-small for calculating embedding-based similarity.
@@
To further improve the model’s ability to leverage retrieved content, we incorporated task-specific RAG examples into the supervised fine-tuning (SFT) stage (1.1.4.b). This strategy—termed Retrieval-Augmented Fine-Tuning (RAFT)—provides the LLM with explicit examples of how to integrate retrieved content during response generation. The evaluation set follows the same construction process as described in Section 1.1.3.a.
@@
We evaluated the impact of RAFT on both in-domain (geoscience) and out-of-domain (general) question-answering tasks. As shown in Table 5, the model trained with RAFT achieved an in-domain recall of 78.12%, compared to 69.72% without RAG-style training—a relative improvement of 12.08%. For out-of-domain tasks, recall increased from 42.04% to 46.49%, a 10.59% gain. These improvements suggest that RAFT enhances the model’s ability to integrate retrieved information into its outputs, leading to more complete, grounded, and contextually accurate responses.
@@ 4 RAG Optimization & Model Fine-tuning
i).SFT Data Synthesis in Geo Science
@@
Instruction: Given the next [document], create a [question] and [answer] pair that are grounded in the main point of the document, don’t add any additional information that is not in the document and 6 The [question] is by an information-seeking user and the [answer] is provided by a helping AI Agent. Refer to the following question format and corresponding answers. Your output should consist solely of question-answer pairs. [question]{few-shot 1 question} [answer] {few-shot 1 answer} …… [document]: contexts ### Response:
@@
In addition to general-purpose datasets, we generate domain-specific data for geoscience using LLaMA3-70B. These synthetic examples are based on abstracts from open-access scholarly articles collected from arXiv. We compile these into a dedicated training set, and apply SimANS to mine hard negatives from the domain corpus, ensuring the model learns to distinguish subtle differences in geoscientific content.
@@
To evaluate in-domain performance, we constructed a dedicated geoscience retrieval benchmark using GeoQA. This benchmark includes approximately 1,000 synthetic query–document pairs generated using LLaMA3-70B and a retrieval corpus of GeoGPT library. As shown in Figure 10, our fine-tuned model achieves substantial gains in retrieval accuracy across all top-$k$ settings—from Recall@1 through Recall@10—demonstrating its effectiveness in domain-specific information retrieval.
@@
Figure 10: Top-$k$ recall on GeoQA benchmark
@@
Dataset and Quality Improvement. To maximize the reranker’s performance in the geoscience domain, we curated a large training corpus composed of both open-source and domain-specific ranking pairs. The domain-specific data were synthesized using LLaMA3-70B, conditioned on abstracts from open-access scholarly articles in geography sourced from arXiv.
@@ 6 Conclusion
In this technical report, we introduced a RAG-based system designed to advance geoscientific research within the GeoGPT framework. The system integrates a complete data processing workflow, robust RAG engineering, a domain-specific embedding model, a domain-specific reranker model, a data synthesis approach, and strategies for enhancing data quality. We conducted comprehensive, multi-level evaluations, all of which demonstrated significant performance improvements. To foster community development and accelerate geoscientific innovation, we have open-sourced the system code, GeoQA dataset, GeoEmbedding model, and GeoReranker model.

### v2

# GeoGPT-RAG Technical Report

@@ 1 Introduction
  • LLM-based synthetic data generation pipeline for geoscience retrieval. To address the lack of high-quality labeled data in the geoscience domain, we design a novel synthetic data generation pipeline. This pipeline produces realistic and diverse training data by abstracting and reformatting scholarly geoscience articles, enabling robust training of retrieval and reranking models under data-scarce conditions.
@@
  • GeoRAG-QA: a new benchmark for geoscience information retrieval. We create and release GeoRAG-QA, the first large-scale benchmark designed for evaluating geoscience retrieval systems. It contains OA paper-based domain-specific documents and 938 expert-reviewed query-document pairs, providing a standardized framework for measuring performance and guiding future research in this area.
@@ 3 RAG Evaluation
We evaluated GeoGPT’s retrieval and generation capabilities using a custom test set generated with the Test Set Generation module from RAGAS (Es et al., 2023). This test set, named GeoRAG-QA, comprises 1,000 questions derived from publicly available geoscience papers. We used gpt-4o as the underlying LLM for both question-answer (QA) pair generation and answer evaluation, and text-embedding-3-small for calculating embedding-based similarity.
@@
To further improve the model’s ability to leverage retrieved content, we incorporated task-specific RAG examples into the supervised fine-tuning (SFT) stage. This strategy—termed Retrieval-Augmented Fine-Tuning (RAFT)—provides the LLM with explicit examples of how to integrate retrieved content during response generation.
@@
We evaluated the impact of RAFT on both in-domain (geoscience) and out-of-domain (general) question-answering tasks. As shown in Table 5, the model trained with RAFT achieved an in-domain recall of 78.12%, compared to 69.72% without RAG-style training—a relative improvement of 12.05%. For out-of-domain tasks, recall increased from 42.04% to 46.49%, a 10.59% gain. These improvements suggest that RAFT enhances the model’s ability to integrate retrieved information into its outputs, leading to more complete, grounded, and contextually accurate responses.
@@ 4 RAG Optimization & Model Fine-tuning
i) SFT Data Synthesis in Geo Science
@@
Instruction: Given the next [document], create a [question] and [answer] pair that are grounded in the main point of the document, don’t add any additional information that is not in the document and [use prompt by different query type of Table 6]. The [question] is by an information-seeking user and the [answer] is provided by a helping AI Agent. Refer to the following question format and corresponding answers. Your output should consist solely of question-answer pairs.
[question]{few-shot 1 question}
[answer] {few-shot 1 answer}
[document]: contexts
### Response:
@@
In addition to general-purpose datasets, we generate domain-specific data for geoscience using Qwen. These synthetic examples are based on abstracts from open-access scholarly articles collected from arXiv under CC BY license, with proper attribution. We compile these into a dedicated training set, and apply SimANS to mine hard negatives from the domain corpus, ensuring the model learns to distinguish subtle differences in geoscientific content.
@@
To evaluate in-domain performance, we constructed a dedicated geoscience retrieval benchmark using GeoRAG-QA. This benchmark includes synthetic query–document pairs generated using Qwen-110B and a retrieval corpus of GeoGPT library. As shown in Figure 10, our fine-tuned model achieves substantial gains in retrieval accuracy across all top-$k$ settings—from Recall@1 through Recall@10—demonstrating its effectiveness in domain-specific information retrieval.
@@
Figure 10: Top-$k$ recall on GeoRAG-QA benchmark
@@
Dataset and Quality Improvement. To maximize the reranker’s performance in the geoscience domain, we curated a large training corpus composed of both open-source and domain-specific ranking pairs. The domain-specific data were synthesized using Qwen, conditioned on abstracts from open-access scholarly articles in geography sourced from arXiv.
@@ 6 Conclusion
In this technical report, we introduced a RAG-based system designed to advance geoscientific research within the GeoGPT framework. The system integrates a complete data processing workflow, robust RAG engineering, a domain-specific embedding model, a domain-specific reranker model, a data synthesis approach, and strategies for enhancing data quality. We conducted comprehensive, multi-level evaluations, all of which demonstrated significant performance improvements. To foster community development and accelerate geoscientific innovation, we have open-sourced the system code, GeoRAG-QA dataset, GeoEmbedding model, and GeoReranker model.